OpenStudy (agentjamesbond007):

Suppose the force of a sliding object is 25 N. Is the force needed to maintain a constant velocity greater than, equal, or less than 25 N?

OpenStudy (kainui):

To maintain a constant velocity, you need a force equal to the force of the object. WHY? Well you know that the unit for force is a Newton. What's a Newton? That's the unit for mass times acceleration, right? F=m*a \[N = kg*m/s ^{2}\] So what's the difference between that and a constant velocity? Well you have an acceleration there _preventing_ you from having a constant speed, right? Afterall, an acceleration is just a change in velocity, right? So you need absolutely no acceleration to have a constant velocity. So if you have a force (mass and acceleration) then you're either speeding up or slowing down. The only way to have a constant speed is to not have a change in speed, and since by definition a force is an acceleration you must have 0 Forces too! So cancel out the 25 N force with another 25 N force in the opposite direction! You will still have momentum and velocity, but you won't be getting faster or slower!